OpenStudy (anonymous):

quantity 1 over 36 minus 1 over x squared all over 1 over 6 plus 1 over x

OpenStudy (anonymous):

Simplify please

OpenStudy (mertsj):

\[\frac{\frac{1}{36}-\frac{1}{x^2}}{\frac{1}{6}+\frac{1}{x}}\]

OpenStudy (mertsj):

Multiply every term of the numerator and denominator by 36x^2. That will cancel all denominators and simplify the complex fraction.

OpenStudy (mertsj):